OpenStudy (anonymous):

Is the derivative of lne^x equal to 0?

OpenStudy (science0229):

No.\[\frac{ d }{ dx }\ln{e^x}=\frac{ d }{ dx }x \ln e=\frac{ d }{ dx }x=1\]

OpenStudy (anonymous):

Thanks!